OpenStudy (anonymous):

have a another question????????? Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (-1, 5) m = -1 is the answer y=-1x+4

OpenStudy (anonymous):

we have : Y2-Y=m(Y2-x) 5-Y=-1(-1-x) 5-y=+1+x -y=x+6 y=-x-6 Got it ?;)

OpenStudy (anonymous):

Look in your equation we have : y=-1 * -1 +4 y=3 is it right ?:)

OpenStudy (anonymous):

Welcome ;)

OpenStudy (anonymous):

i know what u did thankx.....

OpenStudy (anonymous):

;)